# CLiC-it 2024: 10th Italian Conference on Computational Linguistics, Dec 04 — 06, 2024, Pisa, Italy

### v1

@@ Abstract
L’obiettivo di questo lavoro è indagare fino a che punto gli attuali LLM possano apprendere rappresentazioni linguistiche astratte in configurazioni multitask. Abbiamo adottato dati sintetici curati su larga scala di vari problemi BLM in italiano, utilizzandoli per studiarne come le rappresentazioni di frasi costruite da modelli di linguaggio pre-addestrati codificano le informazioni semantiche e sintattiche. Abbiamo utilizzato un’architettura a due livelli per modellare separatamente da un lato la compressione degli embeddings delle frasi in una rappresentazione che contiene informazioni rilevanti per i tasks BLM e dall’altro il BLM stesso. Abbiamo poi verificato se fosse possibile ottenere rappresentazioni compresse di frasi che codificano informazioni sintattiche e semantiche rilevanti per i diversi tasks BLM. Contrariamente alla predizione che la struttura della frase - in termini di sequenza di frasi/chunks - e le proprietà dei chunk possano essere condivise tra i vari tasks, i risultati e l’analisi degli errori mostrano che gli indizi per i diversi task sono codificati in modo diverso negli embeddings delle frasi, suggerendo che nozioni linguistiche astratte come i costituenti o i ruoli tematici non vi sembrano essere presenti.
@@
Blackbird’s Language Matrices (BLM) [2] is a recent task inspired by visual tests of analytic intelligence (RPMs, [3]). The BLM tasks have cast light on whether the correct predictions in previously studied linguistic problems, e.g. number agreement, stem from sentence embeddings that encode deeper linguistic information, such as syntactic structure and semantic properties of phrases [4, 5, 6]. We found that higher-level information – syntactic structure and argument structure – can be assembled from the information encoded in the sentence embeddings. This, however, may not be due to a deeper understanding of such information encoded by LLMs, but rather because of useful surface indicators [7].
@@
We concentrate on Italian, and study several grammatical problems whose solutions can theoretically help each other, in a multi-task setting. We adopt a two-level architecture, developed specifically to model what we know about how humans solve puzzles similar to BLMs [8], to provide insights into how LLMs encode different types of syntactic and semantic information.
We make two contributions: (i) an initial core BLM dataset for Italian that covers linguistic problems of different nature; (ii) single and multi-task experiments that provide new insights into the information encoded by LLMs.
@@ 2 Related Work
We use the multi-learning set-up to integrate tasks across different linguistic levels, syntactic agreement and the syntactic-semantic interface of verb alternation, under the BLM framework. Our aim is to test whether information that can theoretically be shared across these tasks can be distilled from sentence embeddings obtained from a pretrained transformer model. Should that be the case, the joint training would be beneficial because the model would encode shared linguistic features, enhancing its ability to generalise across related linguistic tasks.
@@ 3 The BLM task and the BLM Italian datasets
| Context | ||||
| 1 | NP-sing | PP1-sing | VP-sing | |
| 2 | NP-plur | PP1-sing | VP-plur | |
| 3 | NP-sing | PP1-plur | VP-sing | |
| 4 | NP-plur | PP1-plur | VP-plur | |
| 5 | NP-sing | PP1-sing | PP2-sing | VP-sing |
| 6 | NP-plur | PP1-sing | PP2-sing | VP-plur |
| 7 | NP-sing | PP1-plur | PP2-sing | VP-sing |
| 8 | ??? | |||
| Answer | ||
| 1 | NP-sing PP1-sing et NP2 VP-sing | Coord |
| 2 | NP-plur PP1-plur PP2-sing VP-plur | correct |
| 3 | NP-sing PP-sing VP-sing | WNA |
| 4 | NP-plur PP1-sing PP1-sing VP-plur | WN1 |
| 5 | NP-plur PP1-plur PP2-plur VP-plur | WN2 |
| 6 | NP-sing PP1-sing PP2-sing VP-plur | AEV |
| 7 | NP-sing PP1-plur PP2-sing VP-plur | AEN1 |
| 8 | NP-sing PP1-sing PP2-plur VP-plur | AEN2 |
@@
3.1 BLM-AgrIt – subject-verb agreement in Italian
The BLM-AgrIt dataset is created by manually translating the seed French sentences [4] into Italian by a native speaker, one of the authors. The internal nominal structure in these languages is very similar, so translations are almost parallel. An illustrative, simplified example for Italian is provided in Figure 7, in the appendix. The dataset comprises three subsets of increasing lexical complexity (called Type I, Type II and Type III data below) to test the ability of the system to handle item novelty.
3.2 BLM-CosIt and BLM-OdIt
While BLM-AgrIt tests information about a formal grammatical property, agreement, the Change-of-state (CoS) and Object-drop (OD) alternation datasets test lexical semantic properties of verbs, their ability to enter or not a causative alternation. CoS represents the causative/incohative alternation, where the object of the transitive verb bears the same semantic role (Patient) as the subject of the intransitive verb (L’artista ha aperto la finestra/La finestra si è aperta ‘The artist opened the window’/‘The window opened’). The transitive form of the verb has a causative meaning. In contrast, the subject in OD bears the same semantic role (Agent) in both the transitive and intransitive forms (L’artista dipingeva la finestra/L’artista dipingeva ‘the artist painted the window’/‘the artist painted’) and the verb does not have a causative meaning [17, 18].
BLM-CosIt context and answers
@@ 3.2 BLM-CosIt and BLM-OdIt
The answer set is composed of one correct answer and contrastive wrong answers, all formed by the same four elements: a verb, two nominal constituents and a prepositional phrase. Figure 2 shows the template.111Following [2], we build the errors representing violations of internal (I), external (E) and relational (R) rules of the BLM, and their combination (e.g. IE IER, etc.). This information is used in the first part of the error acronym. The second part of the errors’ label indicates the structure the sentence represent: intransitive (Int), passive (Pass), Transitive (Trans) or, in some cases, the NP introduced by the da preposition (WrBy).
BLM-OdIt Context and Answers
The BLM for OD is the same as for COS, but here the passive voice serves as a confounding element and one of the contrastive answers for CoS is, in fact, the correct answer here.
@@
Levels of abstraction
Each class BLM dataset is developed in two variants, to understand how lexical abstraction impacts learning. One variant (FUN) exhibits functional lexicalisation and includes only elements of the nominal functional lexicon, such as pronominal forms (e.g., Io ’I’, tu ’you’, lei ’she’) and demonstratives (questo ’this’, quello ’that’). The less abstract datasets (LEX) include nominal lexical items of varying complexity. We illustrate the two types of lexical abstraction in Figure 8 in the appendix with the Italian COS verb chiudere ’close’.
@@ 3.3 Dataset statistics
Each subset is split 90:20:10 into train:dev:test subsets. The training and testing are disjoint (agreement data is split based on the correct answer, the alternations data based on the verb). Agreement has 230 test instances for type I, 4121 for types II and III. The verb alternations have 240 test instances for all lexical subsets, for each of the Fun and Lex variations. We randomly sample N training instances, N = 2000.
@@ 4 Multi-task representations
Sentence embeddings encode much information from the input sentence – lexical, syntactic, semantic, and possibly other types of information. Previous experiments have shown that sentence embeddings can be compressed into very small representations (vectors of size 5) that encode information about the structure of the sentence in terms of chunks and their properties, such that they contribute to finding the sequence patterns in BLMs [6]. In this work, we investigate whether several BLM tasks can share the same structural information from a sentence embedding. Towards this end, we built a multi-task version of a two-level system illustrated in Figure 3. In this system, one level processes individual sentences and learns to compress them into small vectors that retain information pertinent to a task, and the other level uses the compressed sentence representation to find patterns across an input sequence to solve a BLM task. The multi-task variation consists in a shared sentence-level, and multiple task components, one for each of the BLM tasks. The system is described in detail in [6]. Both levels are variational encoder-decoders.
Each BLM instance presented to the system is decomposed into individual sentences which are passed through the sentence level. The representation of the input sequence is reassembled from the representations on the latent layer of the sentence level, and passed to the corresponding task module. The input batches are shuffled, to alternate between tasks during training, and avoid getting stuck in a local maximum for one of the tasks.
@@ 5 Multi-task results
Previous published work from our group and current ongoing work has benchmarked the problems generated by some of these datasets [4, 5]. This work has shown that information about the syntactic objects in a sentence and their properties can be obtained from sentence embeddings, and this information is helpful in solving the BLM tasks. We had studied these tasks separately, and investigate here whether such structure is encoded in the sentence embeddings, or whether it is assembled based on shallower patterns within the sentence representations.
@@ Discussion
As the results in Figure 4 show (and also the detailed results in Tables 1-3), single-task training outperforms multi-tasking in the agreement and COS subtasks. The drop suggests that the multi-task model is not able to learn shared properties for these tasks, and forcing it to do so leads to a model that is not optimal for either of them. Both tasks require information about the syntactic structure (or sequence of phrases), while each requires different phrase properties – grammatical number for the agreement task, and semantic properties for the COS. While the system is able to distil all this information from sentence embeddings in the single-task setting, it is not able to compress it into a shared representation when learning the tasks together.
The OD single-task and multi-task have comparable performance, probably because the OD tasks involve a simpler alternation than the COS task. They do not have a causative meaning and do not require a change in the semantic role of the subjects.
@@
For the alternation tasks, error patterns vary, although their distributions remain similar between single task and multi-task environments. We observe an overall increase of error proportions in the multi-task environment. Specifically, mistakes of the type I-int are frequent in type III data for COS Lex. These errors incorrectly map the thematic roles onto the syntax of the arguments (e.g. L’artista si è chiuso ‘the artist closed’ or La carbonara mangiava ‘the carbonara was eating’). In the same dataset, we also note an increase of errors related to the last constituent in type I and type II data (errors of type E-WrBy, e.g. La finestra si chiuse dall’artista ‘the window closed by the artist’). Finally, in OD Lex we remark that R-trans errors are not the most prominent ones, —these are the errors resulting in standard transitive clauses (e.g., L’artista dipinse un paesaggio ‘the artist painted a landscape’)— and do not increase in multi-task environments, suggesting that the chosen answer is not derived from some forms of transitive bias [19].
@@ 6 Conclusions
The results on the performance and the error analysis of a tailored two-level architecture have shown that multi-task environments do not help, or only help only marginally in high-performance settings, suggesting that abstract linguistic notions, such as constituents or thematic roles do not seem to be present in the learning process.
Current work is developing new dedicated architectures based on Variation Autoencoders [5] and creating new BLM problems across various languages and linguistic phenomena.
@@ A.2 Verb alternation examples
| ItaCOSFun - Context | |
| 1 | Io chiudo questa con questo |
| 2 | Io chiudo questa da qui |
| 3 | Questa è chiusa da me con questo |
| 4 | Questa è chiusa da me da qui |
| 5 | Questa è chiusa con questo |
| 6 | Questa è chiusa da qui |
| 7 | Questa si chiude con questo |
| ItaCOSFun - Answers | |
| 1 | Questa si chiude da qui |
| 2 | io mi chiudo da qui |
| 3 | questa è chiusa da me |
| 4 | io sono chiusa da questa |
| 5 | questa chiude me |
| 6 | io chiudo questa |
| 7 | questa si chiude da me |
| 8 | io mi chiudo da questa |
| ItaCOSLex - Context | |
@@ Appendix B Results
| train on | test on | task | ||||
| agreement | COSFun | COSLex | ODFun | ODLex | ||
| type I | type I | 0.904 (0.000) | 0.994 (0.006) | 0.926 (0.006) | 0.992 (0.004) | 0.997 (0.002) |
| type II | 0.772 (0.013) | 0.879 (0.018) | 0.949 (0.020) | 0.971 (0.007) | 0.992 (0.004) | |
| type III | 0.632 (0.019) | 0.949 (0.019) | 0.910 (0.009) | 0.975 (0.007) | 0.918 (0.017) | |
| type II | type I | 0.535 (0.078) | 0.986 (0.006) | 0.893 (0.031) | 1.000 (0.000) | 0.989 (0.006) |
| type II | 0.531 (0.058) | 0.831 (0.010) | 0.971 (0.014) | 0.949 (0.009) | 0.997 (0.002) | |
| type III | 0.438 (0.059) | 0.961 (0.013) | 0.875 (0.014) | 0.996 (0.000) | 0.918 (0.002) | |
| type III | type I | 0.307 (0.009) | 0.992 (0.011) | 0.957 (0.017) | 1.000 (0.000) | 0.990 (0.009) |
| type II | 0.348 (0.002) | 0.900 (0.027) | 0.983 (0.014) | 0.983 (0.007) | 0.999 (0.002) | |
| type III | 0.342 (0.003) | 0.957 (0.010) | 0.871 (0.007) | 0.994 (0.005) | 0.942 (0.018) | |
| train on | test on | |||||
| Fun | Lex | |||||
| type I | type II | type III | type I | type II | type III | |
| type I Fun | 0.999 (0.002) | 0.976 (0.004) | 0.978 (0.009) | 0.151 (0.010) | 0.137 (0.021) | 0.119 (0.012) |
| type II Fun | 0.994 (0.002) | 0.986 (0.004) | 0.987 (0.006) | 0.182 (0.017) | 0.157 (0.016) | 0.139 (0.008) |
| type III Fun | 0.997 (0.002) | 0.987 (0.006) | 0.986 (0.002) | 0.174 (0.007) | 0.126 (0.016) | 0.112 (0.003) |
| type I Lex | 0.306 (0.020) | 0.314 (0.017) | 0.221 (0.030) | 0.632 (0.036) | 0.476 (0.009) | 0.483 (0.032) |
| type II Lex | 0.247 (0.032) | 0.222 (0.014) | 0.175 (0.003) | 0.489 (0.034) | 0.451 (0.019) | 0.549 (0.014) |
| type III Lex | 0.276 (0.012) | 0.236 (0.007) | 0.167 (0.018) | 0.583 (0.010) | 0.517 (0.012) | 0.425 (0.024) |
| train on | test on | |||||
| Fun | Lex | |||||
| type I | type II | type III | type I | type II | type III | |
| type I Fun | 0.993 (0.004) | 0.942 (0.009) | 0.919 (0.009) | 0.193 (0.012) | 0.136 (0.014) | 0.167 (0.016) |
| type II Fun | 0.989 (0.004) | 0.981 (0.005) | 0.956 (0.010) | 0.181 (0.002) | 0.176 (0.007) | 0.151 (0.005) |
| type III Fun | 0.990 (0.002) | 0.996 (0.000) | 0.965 (0.002) | 0.182 (0.004) | 0.164 (0.008) | 0.133 (0.006) |
| type I Lex | 0.369 (0.031) | 0.325 (0.024) | 0.215 (0.012) | 0.728 (0.010) | 0.601 (0.022) | 0.672 (0.025) |
| type II Lex | 0.235 (0.009) | 0.204 (0.010) | 0.165 (0.014) | 0.653 (0.005) | 0.607 (0.024) | 0.681 (0.014) |
| type III Lex | 0.360 (0.009) | 0.314 (0.014) | 0.239 (0.014) | 0.740 (0.004) | 0.678 (0.019) | 0.592 (0.024) |

### v2

@@ Abstract
L’obiettivo di questo lavoro è indagare fino a che punto gli attuali LLM apprendono rappresentazioni linguistiche astratte in configurazioni multitask. Utilizzando dati sintetici curati su larga scala di vari problemi BLM in italiano, studiamo come le rappresentazioni di frasi costruite da modelli di linguaggio pre-addestrati codifichino le informazioni semantiche e sintattiche. Abbiamo utilizzato un’architettura a due livelli per modellare separatamente, da un lato, la compressione degli embeddings delle frasi di input in una rappresentazione che contiene informazioni rilevanti per i tasks BLM e, dall’altro lato, il BLM stesso. Abbiamo poi verificato se fosse possibile ottenere rappresentazioni compresse di frasi che codificano informazioni sintattiche e semantiche rilevanti per i diversi tasks BLM. Contrariamente alla predizione che la struttura della frase - in termini di sequenza di frasi/chunks - e le proprietà dei chunk possano essere condivise tra i vari tasks, i risultati e l’analisi degli errori mostrano che gli indizi per i diversi task sono codificati in modo diverso negli embeddings delle frasi. Questo risultato suggerisce che nozioni linguistiche astratte come i costituenti o i ruoli tematici non vi sembrano essere presenti.
@@
Blackbird’s Language Matrices (BLM) [2] is a recent task inspired by visual tests of analytic intelligence (Raven Progressive Matrices/RPMs, [3]). The BLM tasks have cast light on whether the correct predictions in previously studied linguistic problems, e.g. number agreement or verb alternations, stem from sentence embeddings that encode deeper linguistic information, such as syntactic structure and semantic properties of phrases [4, 5, 6]. We found that higher-level information – syntactic structure and argument structure – can be assembled from the information encoded in the sentence embeddings. This, however, may not be due to a deeper understanding of such information encoded by LLMs, but rather because of useful surface indicators [7].
@@
We concentrate on Italian, and study several grammatical problems whose solutions can theoretically help each other, in a multi-task setting. We adopt a two-level architecture developed specifically to model what we know about how humans solve puzzles similar to BLMs [8]. Level 1 aims to obtain compressed sentence representations that capture information about constituents and their properties; level 2 uses the compressed sentence representations to solve a BLM problem. This architecture provides a tool to study how LLMs encode different types of syntactic and semantic information.
We make two contributions: (i) an initial core BLM dataset for Italian that covers linguistic problems of different nature; (ii) single and multi-task experiments that provide new insights into the information encoded by LLMs. The datasets are available at https://www.idiap.ch/dataset/(blm-agri|blm-causi|blm-odi) and the code at https://github.com/CLCL-Geneva/BLM-SNFDisentangling.
@@ 2 Related Work
Pretrained large language models exhibit general-purpose abilities and knowledge, with high results with little or no fine-tuning on downstream tasks [17, 18]. We can then regard these language models as the results of "multi-task" learning, and our aim here is to test whether sentence embeddings obtained from these models encode syntactic and semantic information consistently, such that different BLM problems that rely on similar linguistic information draw on the same clues from these representations. In particular, we will use BLM tasks on subject-verb agreement – which relies on chunk structure and the chunks’ grammatical number properties – and on verb alternations – which relies on chunk structure and the chunks’ semantic role properties – to test whether chunk structure is encoded in a manner that allows for it to be shared by the two tasks.
@@ 3 The BLM task and the BLM Italian datasets
| BLM agreement problem (BLM-AgrI) | |||
| Context Template | |||
| NP-sg | PP1-sg | VP-sg | |
| NP-pl | PP1-sg | VP-pl | |
| NP-sg | PP1-pl | VP-sg | |
| NP-pl | PP1-pl | VP-pl | |
| NP-sg | PP1-sg | PP2-sg | VP-sg |
| NP-pl | PP1-sg | PP2-sg | VP-pl |
| NP-sg | PP1-pl | PP2-sg | VP-sg |
| Answer set | ||||
| NP-pl | PP1-pl | PP2-sg | VP-pl | Correct |
| NP-pl | PP1-pl | et PP2-sg | VP-pl | Coord |
| NP-pl | PP1-pl | VP-pl | WNA | |
| NP-pl | PP1-sg | PP1-sg | VP-pl | WN1 |
| NP-pl | PP1-pl | PP2-pl | VP-pl | WN2 |
| NP-pl | PP1-pl | PP2-pl | VP-sg | AEV |
| NP-pl | PP1-sg | PP2-pl | VP-sg | AEN1 |
| NP-pl | PP1-pl | PP2-sg | VP-sg | AEN2 |
@@
3.1 BLM-AgrI – subject-verb agreement in Italian
The BLM-AgrI dataset is created by manually translating the seed French sentences [4] into Italian by a native speaker, one of the authors, and then generating the full dataset following the same process of lexical augmentation and sentence shuffling among instances described in [4]. The internal nominal structure in these languages is very similar, so translations are almost parallel. An illustrative, simplified example for Italian is provided in Figure 7, in the appendix. The dataset comprises three subsets of increasing lexical complexity (called Type I, Type II and Type III) to test the ability of the system to handle item novelty.
3.2 BLM-CausI and BLM-OdI
While BLM-AgrI tests information about a formal grammatical property, agreement, the Causative (Caus) and Object-drop (Od) alternation datasets test lexical semantic properties of verbs, their ability to enter or not a causative alternation. Caus represents the causative/inchoative alternation, where the object of the transitive verb bears the same semantic role (Patient) as the subject of the intransitive verb (L’artista ha aperto la finestra/La finestra si è aperta ‘The artist opened the window’/‘The window opened’). The transitive form of the verb has a causative meaning. In contrast, the subject in Od bears the same semantic role (Agent) in both the transitive and intransitive forms (L’artista dipingeva la finestra/L’artista dipingeva ‘the artist painted the window’/‘the artist painted’) and the verb does not have a causative meaning [19, 20].
BLM-CausI context and answers
@@ 3.2 BLM-CausI and BLM-OdI
The answer set is composed of one correct answer and contrastive wrong answers, all formed by the same four elements: a verb, two nominal constituents and a prepositional phrase. Figure 2 shows the template.111Following BLM formal specifications [2], we build the errors representing violations of internal (I), external (E) and relational (R) rules of the BLM, and their combination (e.g. IE IER, etc.). This information is used in the first part of the error acronym. The second part of the errors’ label indicates the structure the sentence represent: intransitive (Int), passive (Pass), Transitive (Trans) or, in some cases, the NP introduced by the da preposition (WrBy).
BLM-OdI Context and Answers
The BLM for Od is the same as for Caus, but here the passive voice serves as a confounding element and one of the contrastive answers for Caus is, in fact, the correct answer here.
@@
We illustrate the data in Figure 8 in the appendix with the Italian Change-of-state verb chiudere ’close’.
@@ 3.3 Dataset statistics
Each subset is split 90:20:10 into train:dev:test subsets. The training and testing are disjoint (agreement data is split based on the correct answer, the alternations data based on the verb). Agreement has 230 test instances for type I, 4121 for types II and III. The verb alternations have 240 test instances for all subsets. We randomly sample a number of training instances, depending on the experimental set-up.
@@ 4 Multi-task representations
Sentence embeddings encode much information from the input sentence – lexical, syntactic, semantic, and possibly other types of information. Previous experiments have shown that sentence embeddings can be compressed into very small representations (vectors of size 5) that encode information about the structure of the sentence in terms of chunks and their properties, such that they contribute to finding the sequence patterns in BLMs [6]. In this work, we investigate whether several BLM tasks can share the same structural information from a sentence embedding. Towards this end, we built a multi-task version of a two-level system, illustrated in Figure 3. In this system, one level processes individual sentences and learns to compress them into small vectors that retain information pertinent to a task and the other level uses the compressed sentence representation to find patterns across an input sequence to solve a BLM task. The multi-task variation consists in a single shared sentence-level component, and multiple task components, one for each of the BLM tasks.
The BLM problems encode a linguistic phenomenon through data that has structure on multiple levels – within sentences, and across a sequence of sentences. We can exploit this structure to develop an indirectly supervised approach to discover and use these different levels of structure. We thus model the solving of a BLM task as a two-step process: (i) compress individual sentences into a representation that emphasizes the sentence structure relevant to the BLM problem (e.g. chunks and their grammatical number for the subject-verb agreement task) (ii) use the compressed representations to detect the sequence-level pattern and solve the BLM task. This two-step process has been shown to be used by people solving visual intelligence tests [21]. In our case, this setup allows us to investigate whether the sentence level can be guided to learn shared information, relevant to the different linguistic tasks described in section 3.
We implement this approach in the two-level intertwined architecture illustrated in Figure 3, and described in detail elsewhere [6]. The data is pre-encoded with Electra [18].222Italian Electra (E-It) pretrained model: dbmdz/electra-base-italian-xxl-cased-discriminator. Multi-lingual Electra (E-M) model: google/electra-base-discriminator. The sentence representations is provided by the embedding of the [CLS] token.333To simplify the discussion of the method, we write ”sentence” instead of ”sentence embedding”, when discussing the system.. We chose Electra because of its stronger sentence-level supervision signal, which leads to higher results when testing the encoding of structural information compared to BERT, RoBERTa, and models tuned by semantic similarity [6].
The two levels are learned together. The input is a BLM instance which is processed on the fly to produce training instances for the sentence level for each sentence in the input sequence . The compressed sentence representations on the latent layer are stacked and passed as input to the task level, which produces a sentence representation as output, which is compared to the answer set of the respective BLM instance .
The sentence level uses a variational encode-decoder architecture to learn how to compress on the latent layer a representation that captures relevant structural information. We guide the system towards this representation by constructing a contrastive set of candidates for comparison with the reconstructed input. The correct output () is the same as the input (), and a selection of other sentences from the input sequence will be the contrastive negative outputs ( (note that an input sequence consists of sentences with different patterns to each other – Figure 1 and 2). We use a max-margin loss function to take advantage of the contrastive answers, is the reconstructed input sentence from the sampled latent vector :
The loss at the task level for input sequence is computed in a similar manner for the constructed answer , but relative to the answer set and the correct answer of the task:
The loss of the two-level systems is:
The input batches are shuffled, to alternate between tasks during training, and avoid getting stuck in a local maximum for one of the tasks.
@@ 5 Multi-task results
Previous published work from our group and current ongoing work has benchmarked the problems generated by some of these datasets [4, 5]. This work has shown that information about the syntactic phrases in a sentence and their properties can be obtained from sentence embeddings, and this information is helpful in solving the BLM tasks. We had studied these tasks separately, and investigate here whether such structure is encoded in the sentence embeddings, or whether it is assembled based on shallower patterns within the sentence representations.
@@ Discussion
As the results in Figure 4 show (and also the detailed results in Tables 1-2 for the Italian Electra pretrained model, and in Tables 3-4 for a multilingual Electra pretrained model), single-task training outperforms multi-tasking in the agreement and verb alternation subtasks. The drop suggests that the multi-task model is not able to learn shared properties for these tasks, and forcing it to do so leads to a model that is not optimal for either of them. Both tasks require information about the syntactic structure (or sequence of phrases), while each requires different phrase properties – grammatical number for the agreement task, and semantic properties for the verb alternation. While the system is able to distil all this information from sentence embeddings in the single-task setting, it is not able to compress it into a shared representation when learning the tasks together.
The Od single-task and multi-task have comparable performance, probably because the Od tasks involve a simpler alternation than the Caus task. They do not have a causative meaning and do not require a change in the semantic role of the subjects.
@@
For the alternation tasks, error patterns vary, although their distributions remain similar between single-task and multi-task environments.
We observe an overall increase of error proportions in the multi-task environment. Specifically, mistakes of the type I-int are frequent in type III data for the Caus task. These errors incorrectly map the thematic roles onto the syntax of the arguments (e.g. L’artista si è chiuso ‘the artist closed’ or La carbonara mangiava ‘the carbonara was eating’). In the same dataset, we also note an increase of errors related to the last constituent in type I and type II data (errors of type E-WrBy, e.g. La finestra si chiuse dall’artista ‘the window closed by the artist’). Finally, for the Od task, we remark that R-trans errors are not the most prominent —these are the errors resulting in standard transitive clauses (e.g., L’artista dipinse un paesaggio ‘the artist painted a landscape’)— and do not increase in multi-task environments, suggesting that the chosen answer is not derived from some forms of transitive bias [22].
@@ 6 Conclusions
The results on the performance and the error analysis of a tailored two-level architecture have shown that multi-task environments do not help, suggesting that abstract linguistic notions, such as constituents or thematic roles do not seem to be present in the learning process.
Current work is developing new analyses and architectures to probe further in the encoding of information in sentence embeddings and creating new BLM problems across various languages and linguistic phenomena.
@@ A.2 Verb alternation examples
| Caus - Context | |
@@ Appendix B Results
B.1 Results with the Italian Electra pretrained model:
dbmdz/electra-base-
italian-xxl-cased-discriminator
| train on | test on | task | ||
| agreement | Caus | Od | ||
| type I | type I | 0.772 (0.011) | 0.910 (0.002) | 0.996 (0.003) |
| type II | 0.660 (0.016) | 0.849 (0.022) | 0.938 (0.007) | |
| type III | 0.483 (0.042) | 0.870 (0.027) | 0.893 (0.010) | |
| type II | type I | 0.504 (0.056) | 0.917 (0.012) | 0.993 (0.004) |
| type II | 0.519 (0.027) | 0.872 (0.007) | 0.981 (0.007) | |
| type III | 0.406 (0.018) | 0.907 (0.004) | 0.950 (0.009) | |
| type III | type I | 0.274 (0.012) | 0.946 (0.003) | 0.994 (0.002) |
| type II | 0.330 (0.004) | 0.929 (0.003) | 0.983 (0.003) | |
| type III | 0.325 (0.008) | 0.889 (0.014) | 0.967 (0.007) | |
| train on | test on | task | ||
| agreement | Caus | Od | ||
| type I | type I | 0.909 (0.007) | 0.919 (0.005) | 1.000 (0.000) |
| type II | 0.760 (0.030) | 0.906 (0.017) | 0.971 (0.003) | |
| type III | 0.707 (0.028) | 0.926 (0.005) | 0.940 (0.010) | |
| type II | type I | 0.881 (0.013) | 0.932 (0.007) | 1.000 (0.000) |
| type II | 0.784 (0.007) | 0.903 (0.010) | 0.983 (0.003) | |
| type III | 0.714 (0.005) | 0.956 (0.005) | 0.975 (0.009) | |
| type III | type I | 0.296 (0.011) | 0.960 (0.005) | 0.998 (0.002) |
| type II | 0.345 (0.002) | 0.950 (0.007) | 0.993 (0.004) | |
| type III | 0.336 (0.005) | 0.918 (0.010) | 0.994 (0.004) | |
B.2 Results with the multilingual Electra pretrained model:
google/electra-base-discriminator
| train on | test on | task | ||
| agreement | Caus | Od | ||
| type I | type I | 0.664 (0.053) | 0.543 (0.011) | 0.714 (0.012) |
| type II | 0.733 (0.018) | 0.407 (0.023) | 0.561 (0.002) | |
| type III | 0.586 (0.022) | 0.483 (0.016) | 0.656 (0.016) | |
| type II | type I | 0.599 (0.025) | 0.610 (0.035) | 0.646 (0.010) |
| type II | 0.660 (0.019) | 0.536 (0.004) | 0.601 (0.004) | |
| type III | 0.518 (0.025) | 0.601 (0.011) | 0.686 (0.019) | |
| type III | type I | 0.320 (0.047) | 0.551 (0.014) | 0.729 (0.015) |
| type II | 0.401 (0.058) | 0.450 (0.021) | 0.661 (0.020) | |
| type III | 0.378 (0.052) | 0.413 (0.012) | 0.618 (0.005) | |
| train on | test on | task | ||
| agreement | Caus | Od | ||
| type I | type I | 0.875 (0.031) | 0.599 (0.040) | 0.749 (0.030) |
| type II | 0.886 (0.005) | 0.425 (0.019) | 0.579 (0.037) | |
| type III | 0.815 (0.016) | 0.529 (0.020) | 0.660 (0.014) | |
| type II | type I | 0.841 (0.024) | 0.543 (0.027) | 0.651 (0.007) |
| type II | 0.881 (0.003) | 0.486 (0.005) | 0.596 (0.010) | |
| type III | 0.814 (0.008) | 0.582 (0.026) | 0.685 (0.013) | |
| type III | type I | 0.826 (0.022) | 0.632 (0.023) | 0.761 (0.023) |
| type II | 0.878 (0.005) | 0.557 (0.013) | 0.697 (0.009) | |
| type III | 0.874 (0.006) | 0.475 (0.010) | 0.592 (0.024) | |